OpenStudy (anonymous):

I need some help please? ill show you the work i did can you please show me where i went wrong?

OpenStudy (anonymous):

sure :)

OpenStudy (anonymous):

\[15\sqrt{10 }* 3\sqrt{8}\] <_____this is the problem i started out with here is the work i did:\[15*3 \] and \[\sqrt{10} * \sqrt{8}\] I end up with: \[45\sqrt{80}\] \[45\sqrt{ 4}\sqrt{20}\] \[45*2\sqrt{20}\] o realized that i would end up with \[90\sqrt{20}\] which is not one of the possible answers....where did i go wrong?

OpenStudy (anonymous):

\[15\sqrt{10} * 3\sqrt{8}\] \[180 \sqrt(5)\]

OpenStudy (anonymous):

Because \[\sqrt{20}\] is the same as \[2\sqrt{5}\] so\[90 \times 2\sqrt{5}\] equals \[180\sqrt{5}\]

OpenStudy (anonymous):

these are the possible answers \[15\sqrt{80}\]\[15\sqrt{18}\]\[45\sqrt{5}\]\[60\sqrt{5}\]

OpenStudy (anonymous):

@some_someone

OpenStudy (anonymous):

maybe you wrote/read, the question wrong; cause that is the corrects answer.